40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 29 13:54:08 2008, rana, Update, SUS, ETMY and PRM have EQ related problems Untitled.png
    Reply  Tue Jul 29 14:38:02 2008, rob, Update, SUS, ETMY and PRM have EQ related problems 
       Reply  Tue Jul 29 19:53:19 2008, Koji, Update, SUS, PRM photos from the south window PRM_from_South_Window1.jpgPRM_from_South_Window2.jpg
Message ID: 756     Entry time: Tue Jul 29 14:38:02 2008     In reply to: 755     Reply to this: 759
Author: rob 
Type: Update 
Category: SUS 
Subject: ETMY and PRM have EQ related problems 

Quote:
The attached trend shows that ETMY and PRM both had large steps in their sensors
around the time of the EQ and didn't return afterwards. The calibration of the
OSEM sensors is ~0.5 mm/V. The PRM sensors respond when we give it huge biases
but there is very little change in the ETMY. Almost certainly true that the
optics have shifted in their wire slings and that we will have to vent to
examine and repair at least ETMY.

Jenne is looking at the spectra of the other suspensions to see if there is
other more subtle issues.


Some additional notes/update:

ETMY, PRM, & MC2 had OSEM signals at a rail (indicating stuck optics). Driving the optics with full scale DAC output freed ETMY and MC2, so while these may have shifted in their slings it may be possible to avoid a repair vent. PRM is still stuck. One OSEM appears to respond with full range to large drives, but the other three face OSEMS remain disturbingly near the rail (HIGH, which is what would happen if a magnet fell off).
ELOG V3.1.3-